OpenStudy (anonymous):

The sum of three consecutive odd numbers is 177. Find the numbers.

OpenStudy (anonymous):

Let the numbers be x , x+2 , x+4 x + x+2 + x+4 = 177 3x + 6 = 177 3x = 171 x = 171 /3 = 57

OpenStudy (anonymous):

so the numbers are 57 , 59 , 61

OpenStudy (ajprincess):

Let the odd numbers be 2x+1,2x+3 and 2x+5. 2x+1+2x+3+2x+5=177 6x+9=177 Find x. then substitute the value of x in 2x+1,2x+3 and 2x+5 to find the numbers

OpenStudy (anonymous):

thnk u